A rectangular surface of length 3.25 cm and width 5.25 cm lies on the x-z plane at y=0. What is the flux through this surface due to a uniform magnetic field of 0.65 T directed 60° from the +x-axis?

1
Expert's answer
2022-05-20T04:53:58-0400
"\Phi=BA\cos\theta\\
=0.65*(0.0325*0.0525)\cos30^\circ=9.6*10^{-4}\:\rm Wb"